OpenStudy (anonymous):

Simplify the expression if possible (4xy^2)^2 and write your answer as a power (using exponents).

OpenStudy (whpalmer4):

\[(4xy^2)^2\] Some handy rules for exponents: \[(ab)^n = a^nb^n\] \[(a^m)^n =a^{m*n}\] That should be all you need. Don't forget that \(x = x^1\)